Cost and Revenues

You have opened your own word-processing service. You bought a personal computer and paid $5,000 for it. However, due to the cost changes in the computer industry, the current price of an equivalent machine is $2,500. You could sell any used machine for $1,000. If you were not word processing, you could earn $20,000 per year at an alternative job. Assume that the interest rate is 10%. You can also hire an assistant who can do everything that you can do for $20,000 per year (you would still continue to do word processing). One person using one computer can produce 11,000 typed pages per year, and the price per page for your service is $2.

 

You are considering three options: (1) expand your business by hiring an assistant, (2) leave your business the way it is, (3) shut down. Based on the costs and revenues above, which should you do? Please explain and show any relevant calculations in more detail.
Answer #1

(A)

Revenue from hiring additional worker = 22,000

Wages = 20,000

Depreciation=1,500 (2500-1000)

Opportunity cost of purchase of new computer = 250 (2500*10%)

Economic profit from additional worker = $250

Total economic profit = (1900+250) $2,150

(B)

Revenue = $22,000

Opportunity cost of your time = 20,000

Opportunity cost of interest on salvage value of existing computer = 100 (1000*10%)

Economic profit = $1,900

You earn $1,900

(C)

Revenue = 0

No costs, since opportunity costs no longer apply,

and fixed costs are sunk.

Economic profit = 0

Option A, expand your business, is the best option.

please hit like you like it.